# Does GenAI Make Usability Testing Obsolete?

### v1

@@ Abstract
Ensuring usability is crucial for the success of mobile apps. Usability issues can compromise user experience and negatively impact the perceived app quality. This paper presents UX-LLM, a novel tool powered by a Large Vision-Language Model that predicts usability issues in iOS apps. To evaluate the performance of UX-LLM we predicted usability issues in two open-source apps of a medium complexity and asked usability experts to assess the predictions. We also performed traditional usability testing and expert review for both apps and compared the results to those of UX-LLM. UX-LLM demonstrated precision ranging from 0.61 and 0.66 and recall between 0.35 and 0.38, indicating its ability to identify valid usability issues, yet failing to capture the majority of issues. Finally, we conducted a focus group with an app development team of a capstone project developing a transit app for visually impaired persons. The focus group expressed positive perceptions of UX-LLM as it identified unknown usability issues in their app. However, they also raised concerns about its integration into the development workflow, suggesting potential improvements. Our results show that UX-LLM cannot fully replace traditional usability evaluation methods but serves as a valuable supplement particularly for small teams with limited resources, to identify issues in less common user paths, due to its ability to inspect the source code.
@@
App Development, Large Language Model, Usability Engineering, AI4SE, Recommender Systems
@@ I Introduction
This work investigates the extent to which GenAI can support or even automate usability evaluation for mobile apps. We introduce UX-LLM, a novel open-source tool that uses Foundation Models to detect usability issues in the individual views of native iOS apps. As input, UX-LLM requires a brief description of the app context, the source code, and an image of the analysed view. Section II introduces the implementation details of UX-LLM and the underlying prompt engineering. This constitutes our first contribution.
To evaluate UX-LLM performance and how it compares with conventional usability evaluation methods, we conducted a multi-method study consisting of expert assessments, expert reviews, and usability testing for two open-source iOS apps: a Quiz and a To-Do app. To understand how development teams perceive the support of such tools and explore possible concerns, we conduced a focus group within an app development project. Section III presents the design of our evaluation study.
@@ II-A Prompt Engineering
Using two own apps previously developed and released on AppStore by the first author, we experimented with various Prompt Engineering strategies and tactics from the OpenAI documentation [^4] https://platform.openai.com/docs/guides/prompt-engineering to enhance the LLM performance. One tactic is to “Ask the model to adopt a persona”, which we used in the system prompt where the LLM is instructed to act as a UX expert. Furthermore, the model was instructed to respond using domain-specific language, avoiding any mention of code. This helps focus the responses, facilitates more user-like feedback, and ensures accessibility to non-technical stakeholders.
@@ II-B Implementation
UX-LLM is a macOS application developed using SwiftUI. When entering an image, it uses TinyPNG to compress it. While waiting for the LLM response, a shader animation is shown, which is sourced from the “Inferno” project by twostraws [^5] https://github.com/twostraws/Inferno. Figure 2 shows a simplified Class diagram of UX-LLM architecture. The diagram particularly abstracts the complexity of the SwiftUI MVVM architecture by condensing the GUI layer and modelling combinations of views and their ViewModels as one. The primary focus is on the service layer, which operates using the “bridge pattern” [23] to separate the abstraction (the high-level logic) from the implementation (the low-level details), allowing them to change independently. This separation enhances modularity and makes the system more flexible and maintainable. The three services PromptGenerator, LLMCaller, and ImageCompressor are abstracted from their actual implementation, allowing UX-LLM to easily adapt to different models or APIs without significant changes. Additionally, the bridge pattern has simplified GUI testing by enabling the substitution of mock objects for actual services.
@@ III Evaluation Design
How accurate can UX-LLM predict usability issues?
@@ III-A Performance Analysis and Methods Comparison (RQ1+2)
For the evaluation of the UX-LLM prediction performance, we first had to select unseen evaluation apps, which we did not use during the development and prompt-engineering. The evaluation apps must have accessible source code that is allowed to be processed by OpenAI’s GPT-4. We weren’t able to use apps from partner companies, as none wanted to share their code with OpenAI due to security and privacy concerns [24]. Consequently, we focused on open-source apps, particularly structurally straightforward apps of medium complexity to ensure a comprehensive, yet manageable evaluation. Furthermore, it was essential to have UX experts assess the predicted usability issues generated by UX-LLM to calculate precision and recall scores, which are established benchmarks for a tool like UX-LLM [25]. In this context, precision calculates of all the issues identified by UX-LLM, how many were actual usability issues. Recall calculates of all the actual usability issues available, how many did UX-LLM successfully identify.
@@
Therefore, we chose a Quiz app [^6] https://github.com/nealarch01/TriviaQuizApp and a To-Do app [^7] https://github.com/fredrik9000/TodoList_SwiftUI shown in Figure 4 and conducted usability evaluations to extract usability issues. The Quiz app consists of four screens. Initially, the user selects a category in which to take a quiz, e.g. Geography. Next, they set up the quiz by adjusting settings, such as the length. Then, they proceed to take the quiz. In the end, there is a score screen that displays the results with options to either retry the quiz or return. The To-Do app consists of two screens. The first screen provides a list of all the tasks that have been added. The second screen is a task detail view that is used to create or edit a task.
@@ III-A2 Expert Review and Expert Assessment
As the second usability evaluation method, we conducted expert reviews with two UI/UX professionals with multiple years of work experience (1 male, 1 female). The first session was conducted in person, and the second via zoom. Each session was divided into two distinct phases, with an introduction at the beginning to explain the obligatory details and the purpose of the study. In the first phase, the experts independently reviewed the two reference apps without knowing about UX-LLM, starting with the Quiz app. They navigated through the interfaces, vocalising their observations and identifying usability issues, similar to the think-aloud protocol. The identified usability issues are listed in the replication package [19].
@@
III-B Perception of Development Teams (RQ3)
@@ IV-A Performance Analysis
View U. Testing Expert Review UX-LLM
Category A1 B4 /
View A2 B5 /
(Quiz App) A3 / C6
/ B2 C1, C12, C20, C23
/ B6 C4
Setup View A5 B9, B10, B11, B14 C8, C14
(Quiz App) A6 B13 /
A7 / C11
Quiz View A9, A10 B21 C15, C21
(Quiz App) A14 B20 C22
/ B18 C19
Score View A15, A16 B29 C25
(Quiz App) / B23, B26 C24
List View A17 B35 /
(To-Do App) A18 B37 C37
A19 B36 /
A21 / C33
A22 B43 C35
/ B38 C32
/ B39 C36
Task Detail A25 B57 C44, C47
View A26 B54 C46
(To-Do App) A27 B58 /
/ B44 C42
/ B51 C41
/ B53 C43
@@ IV-B Comparison of Usability Evaluation Methods
IV-C Perception of Development Teams (RQ3)
@@ IV-C4 Additional concerns and improvement suggestions
For applying UX-LLM in daily work, the group proposed embedding it as a plugin in the IDE or in a continuous integration (CI) pipeline to allow for automatic operation in the background. They emphasised that an IDE plugin “just makes sense, as the Xcode IDE already has the code and view’s preview side by side”. In addition, the team expressed the desire to not only identify issues but also provide solutions, including alternative designs of their input image. The designer said: “When it criticised the accessibility of the colours, it would be nice if it could also show what colours to use instead”. Interestingly, they also stated a limitation that could result from the policy of the LLM used. A discussed example is when creating an app about wine, as some LLMs might not be allowed to answer queries about alcohol. Furthermore, the presentation of usability issues was relevant to the team, as they found it demotivating to receive a lengthy list pointing out their product deficiencies. Also, the developer was sceptical about UX-LLM capability to identify usability issues in features reliant on hardware interactions involving user input or external factors, such as camera functions and voice input. Lastly, they emphasised the importance of a holistic analysis to detect broader inconsistencies and navigation issues. This approach would also reduce predicted usability issues, which are already addressed in other sections of the app.
@@ V-2 Usage Data Analytics
App usage logs deliver great value for developers as they enable analysing the user behaviour [48, 49, 50]. For example, they can identify which app features are used most and focus usability evaluation on those areas [51, 52]. Mobile tracking [53, 52] aka analytics tools like Firebase or TelemetryDeck enable exploring runtime, engagement, and navigation data, or tracking actions such as button presses. However, such tracking remains exploratory and raises privacy concerns depending on its implementation [54, 24].
@@ V-3 User Feedback Analysis
Research has shown that, for many developers, App Store reviews are a primary source of user feedback [57, 58]. These reviews enable gathering issues including on usability and help developers improve their apps, enabling user-driven quality assessment [48]. Users also report issues in other feedback channels too as social media and product forums. Nayebi and Cho [59], e.g., revealed that more than 12% of bug reports were discovered on Twitternow called “X”. Thus, monitoring and processing feedback looking for usability related reports is an active field of research that received a boost with the availability of recent LLMs [60]. While certainly an important complementary source for monitoring usability issues, actual user priorities, and novel ideas, user feedback analytics also requires releasing the app and having a significant user base—whereas UX-LLM can be used during the development as it simulates common (user) knowledge coded in foundation models. UX-LLM is also beneficial for less popular apps, which do not have enough users for a meaningful analysis [58]. Feedback analytics also requires manual tuning and interpretation to extract more valuable insights [61, 62]. Feedback can vary greatly in quality with up to 75% of app store reviews consist of praise rather than constructive feedback [58, 35].
@@ V-4 Generative AI for GUI Engineering
Recent advance in AI has also benefited usability research and UI design in general. Beltramelli [63] presented “pix2code”, which generates code from a GUI image input. This method uses Convolutional Neural Networks to identify distinct elements, positions, and poses in GUI images. It then employs Recurrent Neural Networks and a decoder to generate executable code. With the rapid raise of Generative AI significant improvements in how models solve the “Design2Code” task. Si et al. [64]. conduct benchmarking with current state-of-the-art models and showed that GPT-4 Turbo with Vision performs best on this task, compared to other models like Gemini Pro. Their study annotators claimed that GPT-4 generated websites which can replace the original reference in terms of visual appearance and content in 49% of cases. In 64% of cases, the annotators stated that GPT-4 even generated better websites than the original reference. Lee et al. [65] proposed a hybrid neural design network to generate design layouts that meet user-specified constraints (e.g. a design with one headline and two images). In their experiments, they demonstrate that the generated layouts are visually similar to real design layouts. Wei et al. [66] take a step further to generate whole GUI designs using UI-Diffuser and LayoutDM [67], a Transformer-based model to generate layouts. Users provide a simple prompt of what the GUI should represent, e.g., a music player app. The preliminary results indicate that this could be an efficient and cost-effective method for creating mobile GUI designs, but far from directly reusable and need improvement and usability evaluation. is a similar recent commercial product, which generates mobile GUI designs with a simple prompt. Mattisson et al. [68] demonstrated that Uizard can significantly enhance designers efficiency, boost creativity in idea generation, and improve communication with stakeholders. Particularly the editable designs with nested components allow for continuous customisation and adjustments, e.g. after usability evaluation. To the best of our knowledge, this work is the first to use GenAI and empirically study its potential for usability evaluation.
@@ VI-1 Internal and External Validity
The selection of reference apps restricts the broader unreliability of our findings. The chosen Quiz and To-Do app are simpler compared to the wide range of available more complex apps. To increase the generalisability, it is important to replicate our study with apps from various categories, embodying varying levels of complexity and incorporating different types of user interactions [50, 71]. The amount and quality of code entered into UX-LLM can also have a major impact on performance.
@@ VI-2 Construct Validity
During the expert reviews, the survey had four options to categorise UX-LLM’s usability issues: “Usability Issue”, “No Usability Issue”, “Uncertain”, and “Incorrect/Irrelevant Statement”. The appropriateness of these categories could also be questioned, considering that they might not capture all possible reactions. To mitigate this potential threat, we took detailed notes of all expert comments.

### v2

@@ Abstract
Ensuring usability is crucial for the success of mobile apps. Usability issues can compromise user experience and negatively impact the perceived app quality. This paper presents UX-LLM, a novel tool powered by a Large Vision-Language Model that predicts usability issues in iOS apps. To evaluate the performance of UX-LLM, we predicted usability issues in two open-source apps of a medium complexity and asked two usability experts to assess the predictions. We also performed traditional usability testing and expert review for both apps and compared the results to those of UX-LLM. UX-LLM demonstrated precision ranging from 0.61 and 0.66 and recall between 0.35 and 0.38, indicating its ability to identify valid usability issues, yet failing to capture the majority of issues. Finally, we conducted a focus group with an app development team of a capstone project developing a transit app for visually impaired persons. The focus group expressed positive perceptions of UX-LLM as it identified unknown usability issues in their app. However, they also raised concerns about its integration into the development workflow, suggesting potential improvements. Our results show that UX-LLM cannot fully replace traditional usability evaluation methods but serves as a valuable supplement particularly for small teams with limited resources, to identify issues in less common user paths, due to its ability to inspect the source code.
@@
App Development, Large Language Model, Foundation Models, Usability Engineering, AI4SE, Recommender Systems, Quality Requirements, AI-Inspired Design.
@@ I Introduction
This work investigates the extent to which GenAI can support or even automate usability evaluation for mobile apps. We introduce UX-LLM, a novel open-source tool that uses Foundation Models to detect usability issues in the separate views of native iOS apps. As input, UX-LLM requires a brief description of the app context, the source code, and an image of the analysed view. Section II introduces the implementation details of UX-LLM and the underlying prompt engineering. This constitutes our first contribution.
To evaluate UX-LLM performance and how it compares with conventional usability evaluation methods, we conducted a multi-method study consisting of expert assessments, expert reviews, and usability testing for two open-source iOS apps of medium complexity: a Quiz and a To-Do app. To understand how development teams perceive the support of such tools and explore possible concerns, we conduced a focus group within an app development project. Section III presents the design of our evaluation study.
@@ II-A Prompt Engineering
Using several app projects which we had access to and were familiar with (but different from those used in the evaluation), we experimented with various Prompt Engineering strategies and tactics from the OpenAI documentation [^4] https://platform.openai.com/docs/guides/prompt-engineering to enhance the LLM performance. One tactic is to “Ask the model to adopt a persona”, which we used in the system prompt where the LLM is instructed to act as a UX expert. Furthermore, the model was instructed to respond using domain-specific language, avoiding any mention of code. This helps focus the responses, facilitates more user-like feedback, and ensures accessibility to non-technical stakeholders.
@@ II-B Implementation
UX-LLM is a macOS application developed using SwiftUI. When entering an image, it uses TinyPNG API [^5] https://tinypng.com to compress it. While waiting for the LLM response, a shader animation is shown, which is sourced from the Inferno project by twostraws [^6] https://github.com/twostraws/Inferno. Figure 2 shows a simplified Class diagram of UX-LLM architecture. The diagram particularly abstracts the complexity of the SwiftUI MVVM architecture by condensing the GUI layer and modelling combinations of views and their ViewModels as one. The primary focus is on the service layer, which operates using the “bridge pattern” [23] to separate the abstraction (the high-level logic) from the implementation (the low-level details), allowing them to change independently. This separation enhances modularity and makes the system more flexible and maintainable. The three services PromptGenerator, LLMCaller, and ImageCompressor are abstracted from their actual implementation, allowing UX-LLM to easily adapt to different models or APIs without significant changes. Additionally, the bridge pattern has simplified GUI testing by enabling the substitution of mock objects for actual services.
@@ III Evaluation Design
How accurately can UX-LLM predict usability issues?
@@ III-A Performance Analysis and Methods Comparison (RQ1+2)
For the evaluation of the UX-LLM prediction performance, we first had to select evaluation apps, which we did not use during the development and prompt-engineering. The evaluation apps must have accessible source code that is allowed to be processed by OpenAI’s GPT-4. We weren’t able to use apps from partner companies, as none wanted to share their code with OpenAI due to security and privacy concerns. Consequently, we focused on open-source apps, particularly structurally straightforward apps of medium complexity to ensure a comprehensive, yet manageable evaluation. Furthermore, it was essential to have UX experts assess the predicted usability issues generated by UX-LLM to calculate precision and recall scores, which are established benchmarks for a tool like UX-LLM [24]. In this context, precision calculates of all the issues identified by UX-LLM, how many were actual usability issues. Recall calculates of all the actual usability issues available, how many did UX-LLM successfully identify.
@@
We followed a systematic approach for the app selection. We first searched GitHub for “SwiftUI” and Google for “Open Source + SwiftUI”. We then manually checked 50 apps looking for structurally straightforward apps of medium complexity to ensure a comprehensive, yet manageable evaluation. We excluded niche, very specialized, or inactive apps. Finally we chose two app: a Quiz app [^7] https://github.com/nealarch01/TriviaQuizApp and a To-Do app [^8] https://github.com/fredrik9000/TodoList_SwiftUI shown in Figure 4. The Quiz app consists of four screens. Initially, the user selects a category in which to take a quiz, e.g. Geography. Next, they set up the quiz by adjusting settings, such as the length. Then, they proceed to take the quiz. In the end, there is a score screen that displays the results with options to either retry the quiz or return. The To-Do app consists of two screens. The first screen provides a list of all the tasks that have been added. The second screen is a task detail view that is used to create or edit a task.
@@ III-A2 Expert Review and Expert Assessment
As the second usability evaluation method, we conducted expert reviews with two UI/UX professionals (1 male, 1 female) each with six years of work experience. Their daily work is to conduct usability evaluations and UX design. The first session was conducted in person, and the second via zoom. Each session was divided into two distinct phases, with an introduction at the beginning to explain the obligatory details and the purpose of the study. In the first phase, the experts independently reviewed the two reference apps without knowing about UX-LLM, starting with the Quiz app. They navigated through the interfaces, vocalising their observations and identifying usability issues, similar to the think-aloud protocol. The identified usability issues are listed in the replication package [19].
@@
III-B Perception of a Development Team (RQ3)
@@ IV-A Performance Analysis
View U. Testing Expert 1 Expert 2 UX-LLM
Category A1 B4 B4 /
View A2 B5 / /
(Quiz App) A3 / / C6
/ / B2 C1, C12, C20, C23
/ / B6 C4
Setup View A5 B10, B11, B14 B9 C8, C14
(Quiz App) A6 / B13 /
A7 / / C11
Quiz View A9, A10 B21 / C15, C21
(Quiz App) A14 B20 / C22
/ B18 B18 C19
Score View A15, A16 B29 / C25
(Quiz App) / B23, B26 / C24
List View A17 / B35 /
(To-Do App) A18 B37 B37 C37
A19 B36 / /
A21 / / C33
A22 / B43 C35
/ B38 / C32
/ B39 / C36
Task Detail A25 B57 / C44, C47
View A26 B54 / C46
(To-Do App) A27 B58 / /
/ B44 / C42
/ B51 B51 C41
/ B53 / C43
@@ IV-B Comparison of Usability Evaluation Methods
IV-C Perception of a Development Team (RQ3)
@@ IV-C4 Additional concerns and improvement suggestions
For applying UX-LLM in daily work, the group proposed embedding it as a plugin in the IDE or in a continuous integration (CI) pipeline to allow for automatic operation in the background. They emphasised that an IDE plugin “just makes sense, as the Xcode IDE [^9] https://developer.apple.com/xcode/ already has the code and view’s preview side by side”. In addition, the team expressed the desire to not only identify issues but also provide solutions, including alternative designs of their input image. The designer said: “When it criticised the accessibility of the colours, it would be nice if it could also show what colours to use instead”. Interestingly, they also stated a limitation that could result from the policy of the LLM used. A discussed example is when creating an app about wine, as some LLMs might not be allowed to answer queries about alcohol. Furthermore, the presentation of usability issues was relevant to the team, as they found it demotivating to receive a lengthy list pointing out their product deficiencies. Also, the developer was sceptical about UX-LLM capability to identify usability issues in features reliant on hardware interactions involving user input or external factors, such as camera functions and voice input. Lastly, they emphasised the importance of a holistic analysis to detect broader inconsistencies and navigation issues. This approach would also reduce predicted usability issues, which are already addressed in other sections of the app.
@@ V-2 Usage Data Analytics
App usage logs deliver great value for developers as they enable analysing the user behaviour [47, 48, 49]. For example, they can identify which app features are used most and focus usability evaluation on those areas [50, 51]. Mobile tracking [52, 51] aka analytics tools like Firebase  [^10] https://firebase.google.com or TelemetryDeck  [^11] https://telemetrydeck.com enable exploring runtime, engagement, and navigation data, or tracking actions such as button presses. However, such tracking remains exploratory and raises privacy concerns depending on its implementation [53, 54].
@@ V-3 User Feedback Analysis
Research has shown that, for many developers, App Store reviews represent a primary source of user feedback [57, 58]. These reviews enable gathering issues (including on usability) and help developers improve their apps, enabling user-driven quality assessment [47]. Users also report issues in other feedback channels too as social media [34] and product forums [59]. Thus, monitoring and processing feedback looking for usability related reports is an active research field that has received a strong boost with the availability of recent LLMs [60, 61]. While certainly an important complementary source for monitoring usability issues, actual user priorities, and novel ideas, user feedback analytics also requires releasing the app and having a significant user base—whereas UX-LLM can be used during the development as it simulates common (user) knowledge coded in foundation models. UX-LLM is also beneficial for less popular apps, which do not have enough users for a meaningful analysis [58]. Moreover, feedback analytics requires manual tuning and interpretation to extract more valuable insights [62, 63, 60]. Feedback can vary greatly in quality [60] with up to 75% of app store reviews consist of praise rather than constructive feedback [58, 34].
@@ V-4 Generative AI for GUI Engineering
Recent advance in AI has also benefited usability research and UI design in general. Beltramelli [64] presented “pix2code”, which generates code from a GUI image input. This approach uses Convolutional Neural Networks to identify distinct elements, positions, and poses in GUI images. It then employs Recurrent Neural Networks and a decoder to generate executable code. The rapid raise of Generative AI brought significant improvements in how models solve the “Design2Code” task. Si et al. [65]. conducted a benchmarking with state-of-the-art models and showed that GPT-4 Turbo with Vision performs best on this task, compared to other models like Gemini Pro. Their study annotators claimed that GPT-4 generated websites which can replace the original reference in terms of visual appearance and content in 49% of cases. In 64% of cases, the annotators stated that GPT-4 even generated better websites than the original reference.
Other recent approaches has suggested to generate GUI images. Lee et al. [66] proposed a hybrid neural network to generate design layouts that meet user-specified constraints (e.g., a design with one headline and two images). In their experiments, they demonstrated that the generated layouts are visually similar to real design layouts. Wei et al. [67] took a step further to generate the designs using UI-Diffuser and LayoutDM [68], a Transformer-based model for layouts generation. Users provide a simple prompt of what the GUI should represent, e.g., a music player. Their preliminary results indicate that this could be a cost-effective approach for creating mobile GUI designs. However, the GUIs are far from being directly directly reusable, need improvements, and usability evaluation. Uizard [^12] https://uizard.io is a similar recent commercial product, which generates mobile GUI designs with a simple prompt. Mattisson et al. [69] demonstrated that Uizard can significantly enhance designers efficiency, boost creativity in idea generation, and improve communication with stakeholders. Particularly the editable designs with nested components allow for continuous customisation and adjustments, e.g. after usability evaluation. In another recent study, Wei et al. [70] tuned a Vision-Language Model to retrieve relevant GUI designs from a large constructed dataset. This approach seems to be more effective yielding more relevant and higher quality screens—highlighting the importance of the human-in-the-loop for GUI recommendation.
The works are are perhaps closest to ours are of Kuang et al. [71] and of Kocaballi [72]. To explore how UX evaluators use AI assistant Kuang et al.[71] conducted a study with 20 participants using a simulated AI assistants. They found that participants asked the bot for five categories of information such as user actions and mental model and observed other trends. The authors finally derive design considerations for future conversational AI assistants for UX evaluation. In hypothetical design project, Kocaballi used ChatGPT to generate personas, simulate interviews with fictional users, and create new design ideas. The work highlights drawbacks such as forgotten information, partial responses, and a lack of output diversity, suggesting the importance of human oversight in the design process. To the best of our knowledge, this work is the first to use Generative AI for predicting actual usability issues in apps while comparing the performance with conventional usability testing and evaluation by humans.
@@ VI-1 Internal and External Validity
The selection of two reference apps restricts a broader reliability of our findings. The chosen Quiz and To-Do app are simpler compared to the wide range of more complex available apps. As we primarily aimed to support smaller app teams, it made more sense to focus the evaluation on simpler apps. Evaluating UX-LLM for niche domains like “Ireland - COVID Tracker” would rather limit the results. The two apps used in our study provided the best compromise between:
  • Broadly used and maintained app.
  • Clean SwiftUI Model-View-ViewModel Architecture to inject into UX-LLM.
  • General apps, not too niche domain (also to enable the usability testing).
. Nevertheless, to increase the generalisability, it is important to replicate our study with apps from various categories, embodying varying levels of complexity and incorporating different types of user interactions [49, 75]. The amount and quality of code entered into UX-LLM can also have a major impact on its performance.
@@ VI-2 Construct Validity
Unlike the usability testing and the expert assessment which followed strict protocols, during the expert reviews we refrained from instructing the experts to follow a certain process. A more structured review would likely increase the replicability and agreement rate. In fact, we considered using the more structured approach called “Heuristic Expert Reviews” where evaluators are aided by a checklist. However, we decided for the general review form: mainly to not interfere or direct the experts to a certain direction [77]. Moreover, heuristic evaluations tend to miss context-specific issues [77].
During the expert assessment, the survey had four options to categorise UX-LLM’s usability issues: “Usability Issue”, “No Usability Issue”, “Uncertain”, and “Incorrect/Irrelevant Statement”. The appropriateness of these categories could also be questioned, considering that they might not capture all possible reactions. To mitigate this potential threat, we took detailed notes of all expert comments.